Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Friday, January 08, 1999

The following is a brief summary of the environmental monitoring activities that were completed between January 08, 1999 1:00 and 1:00 on January 9, 1999. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

01/08/99 (Friday)

Segment 1 (Emidio Pump Station to Copco Road): Report for 8 January 1999: Fiber optic crews worked on a ~9000' section of cable today, between Tejon Ranch and south of Cuddy Creek. Backfill progressed at the Warbler Skip, with the crew completing ~1300 feet (of 1600'). Clean up of this skip also began today. Work on Tejon Bench consisted of continuing backfill and compaction, as well as minor clean up. Crews on the west side of the I-5/Tejon bore spent the day digging around foreign lines. A shoring box was removed, with much difficulty, from the exit hole for the Grapevine Creek bore. Crews began digging the rest of the ditch out, while crews on the north side worked on the tie in. Paving, as well as clean up work, was done on the I-5 frontage road. Welding and x-ray were performed on ~1500 feet of surplus pipe strung along Highway 166. No work occurred at Emidio Station, while crews at Grapevine Station worked on gauges.

Segment 2B (Castaic to Sylmar). Report for 01-08-99. Electrical work continued at the Church street vault site. Valve and electrical work occurred at the Live Oak, Castaic Creek, Balboa vaults and the vault north of Calgrove Rd. clean-up and electrical work took place near Weldon Canyon road.

Segment 3 (Sylmar to Los Angeles). Report for 01-8-99. Work continued at the Tie-in near Balboa Blvd. The Vault at Balboa was the site of valve and electrical work. A street capping crew worked on a section of San Fernando Road, south of Balboa. A rubber -tired hoe worked on trenching for electrical conduit at Highland Ave, where a truck delivered slurry. A crew continued to work on the Los Feliz Bridge crossing.

Segment 5 (Lynwood to Long Beach): Report for 1-8-99. Crews progressed with the brickwork and began hanging doors at Lynwood station. Meanwhile near the 91 Fwy, crews continued clean up and grading of the rail right of way. Work at Edison continued with boring crews setting the rails and initiating the bore under Alameda St. In addition, crews extended the traffic control and placed flaggers along the roadway to insure a safe work area. Electrical crews worked setting boxes at Tosco, backfilling and grading the right of way. Likewise, at Texaco crews backfilled and cleaned up the general work area. Segment 4B (Lynwood to El Segundo): Crews cleaned out dirt from the paved areas in anticipation for concrete and asphalt to be poured. In addition, electrical crews continued final work on the skids and excess spoils were removed from the work area.

Inspection Summar

In the past 24 hours, 4 inspection event(s) were reported for 4 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page